### v1

@@ I Introduction
Thermodynamics is a remarkable theory in physics, primarily aimed at studying heat [1]. It has been instrumental in the design of internal combustion engines and even in explaining the behavior of black holes. Technological progress increasingly requires the miniaturization of components [2], pushing them to a regime where quantum effects are no longer negligible. The necessity of understanding the underlying mechanisms of converting heat into useful work has led researchers to formulate the three laws of thermodynamics. More recently, with the rise of quantum technologies and the miniaturization of devices that exchange heat and work at the nanoscale, it has become increasingly relevant to investigate these mechanisms within the framework of quantum mechanics [3]. However, while classical thermodynamics is a well-established theory, its extension to the quantum domain presents conceptual challenges and remains an active area of research. Pioneering efforts to generalize the concepts of heat and work for quantum systems date back to the 1980s [4]. Only recently has a renewed interest in quantum thermal machines spurred prolific scientific advancements, starting with the groundbreaking proposal of the maser [5] as the first example of a quantum machine.
@@
A quantum thermal machine is broadly defined as a device made of quantum systems capable of performing work by undergoing a suitable thermodynamic cycle. Several studies have proposed potential implementations of few-body thermal machines [6] based on the Carnot [8, 7], Otto [9, 10], Stirling [11, 12], Brayton [13], and Diesel cycles [14]. Some of these realizations have been achieved using various quantum systems, including infinite potential wells [15], nuclear magnetic resonance (NMR) [16], ultra-cold atoms [17]. However, the role played by many-body interactions in the thermodynamic performance of quantum machines [18, 19, 20] is still not fully understood. Determining whether quantum machines with many-body interactions can outperform classical ones, and under what conditions this occurs, is currently at the forefront of debates in quantum thermodynamics. Long-range interacting systems [21, 22] offer promising prospects for quantum technological applications due to their resilience to external disturbances [23]. This stability enables control over the impact of dynamically generated excitation, thereby mitigating their detrimental effects. Specifically, the dynamical phase transition [24, 25] is clear when $\alpha<1$. But when $\alpha>1$, the critical point grows into a chaotic crossover region where the dynamics and asymptotic state are very dependent on the system’s parameters.
A promising direction for exploring quantum advantage lies in using many-body quantum systems as the working substance in thermodynamic engines [26, 27]. Although this approach is hindered by technical challenges, both analytically and numerically, some preliminary results have been obtained [28, 29, 30]. It has been shown that quantum criticality can play a crucial role in enhancing engine performance [31, 32, 33]. However, the exact role that many-body interactions play in the thermodynamic efficiency of quantum engines remains poorly understood. Besides, determining whether quantum engines featuring such interactions can outperform classical ones, and under what conditions, is a key question in the current quantum thermodynamics debate. One of the main obstacles lies in the difficulty of solving the dynamics of these many-body systems. In this paper, we contribute to this burgeoning field by presenting an in-depth study of the quantum Otto engine based on the Ising model (see Fig.1), focusing particularly on the potential influence of quantum criticality on the engine’s thermodynamic performance.
This article is structured as follows: First, we will present our system and the Otto cycle, describing its applications. We will explore the impact of internal parameters such as $N$, $h_{i}$, $\alpha$, $\beta_{c}$, and $\beta_{h}$ on the operating modes and highlight the quantum features absent in classical systems, as well as the transition between them. Next, we will look at the quantum refrigerator (or thermal engine) mode and see how the chain size (N) and thermal correlations change the quantum refrigerator capability $Q_{c}/N$ (or $W/N$) and the efficiency $\eta_{R}$ (or $\eta$). We will fix the time of adiabatic compression throughout the study, unlike in previous works. We will address the scenario of a perfect quantum thermodynamic machine and a refrigerator, both of which experience an infinitely slow thermodynamic phase. We will calculate the work per spin, the efficiency, and the influence of $N$ on the extraction of work per spin and its effect on efficiency. We will also perform the same analysis for the refrigerator and also describe the performance and the influence of perturbations and dissipation, using the scaling factor [34] $\Pi_{R}/N$ (or $\Pi/N$), which serves as a reliable indicator for stability. In the following section, we will examine the behavior of correlations [24, 25] for $\alpha$ ranging from 0.01 to 1.3 with a fixed $N=100$, focusing on quantum refrigerators and thermal engines. We will study how the efficiency behaves as a function of $Q_{c}/N$ and $h_{i}$ (or $\eta$ as a function of $W/N$ and $h_{i}$), and analyze their distribution across chains of varying sizes. Our final step is to calculate the proportionality between the scaling factor per spin for $N$ from $10$ to $100$, as well as the behavior of peaks before and after the phase transition and at the critical field.
@@ II.1 Many-body quantum Ising model
ensures that the energy density remains intensive for $\alpha\leq 1$. For $\alpha=\infty$, this model reduces to the standard transverse field Ising model (TFIM), which can be exactly solved using the Jordan-Wigner transformation [36, 35] . The long range exhibits a quantum phase transition from the ferromagnetic to the paramagnetic phase at $h=2J$. This transition persists as $\alpha$ decreases [37], with the transition point shifting towards higher magnetic field values.
@@
At the opposite extreme, $\alpha=0$ [38, 39]represents a fully connected regime, which can be analytically treated for its static and dynamic properties. In this case, the model shows long-range ferromagnetic order at low finite temperatures. This configuration is particularly interesting for $\alpha<2$, revealing exotic phenomena such as prethermalization [21, 40], dynamical phase transitions [25, 24], and dynamical confinement [41, 42].
@@
This model is strongly influenced by the parameter $\alpha$. For instance, for $\alpha<2$, long-range ferromagnetic order is observed at finite temperatures [43], along with unique dynamic behaviors that have garnered significant experimental interest. Trapped ion systems, with their adjustable transverse fields and long-range interactions, provide an ideal framework for studying these phenomena. The critical magnetic field is estimated using the Cluster approach and Mean-Field Theory (CMFT) [44, 45, 46] to explore the phase diagram of an Ising spin chain with long-range interactions. This method captures short-range correlations while treating interactions between clusters at a mean-field level, revealing the emergence of a chaotic region for specific values of the parameter $\alpha$. In our study, we avoid mean-field or cluster decomposition approximations and spin interaction is uniform. Nevertheless, we observe significant concordance with the results obtained via CMFT, especially concerning static phase transitions. This agreement underscores the robustness of our method, providing an accurate description of phase transitions without resorting to simplifications. From this observation, it is possible to extract a qualitative behavior of the quantum critical point as a function of $\alpha$. Interpolating the data, we find
@@
Since our method is essentially a generalized mean field theory, we will use it when the latter describes the equilibrium transition ($\alpha<5/3$). Despite this, it must be stressed that CMFT is exact for large $\ell$. We will use a numerical diagonalization of the Hamiltonian in the disordered case similar to [10], where we will assume $J_{ij}=\frac{1}{K(\alpha)}\frac{J}{|i-j|}$, to compute $\omega_{k}(h)$ for more details consult Otto [35]. In conclusion, the transverse field Ising model with power-law decaying interactions offers a rich theoretical framework for exploring quantum phase transitions and complex dynamics while being directly applicable to concrete experimental setups.
@@ II.2 Cycle description
It describes a one-dimensional long-range Ising system of $N$ quantum spins interacting with a ferromagnetic coupling strength $J>0$, in the presence of magnetic field $h$. Hereafter, we set $J=0,5$ as the energy scale and work in units where $\hbar=k_{B}=1$.
The long-range interaction between spins is dictated by the strength parameter $\alpha$ (we set $h_{i}<h_{f}$ and $h_{f}-h_{i}=0.5$). We will work on an ideal otto cycle illustrated in Fig.1 composed of two infinitely slow adiabatic phases and two thermalization phases without any dissipation. After that, we will examine the influence of the thermal correlations of the $N$ particles on the efficiency of the cycle. For the dynamic and ideal case, we have:
A → B: Adiabatic increase of the transverse field
The field $h(t)$ increases linearly from $h_{i}$ to $h_{f}$, while the working substance is kept isolated from the thermal baths. The expression for $h(t)$ is:
$h(t)=h_{i}+0.5\frac{t}{T}$
  • For the ideal case: $T\to\infty$ and $h=h_{i}$
B → C: Thermalization with the cold bath
Once the field reaches $h_{f}$, the magnetic field is fixed, and the working substance is brought into contact with the cold bath until it reaches thermal equilibrium at temperature $T_{c}$ (we will assume a perfect thermalization).
C → D: Adiabatic decrease of the transverse field
The field $h(t)$ is decreased linearly from $h_{f}$ back to $h_{i}$ at the same rate as in step A → B, while keeping the working substance isolated from the thermal baths. The expression for $h(t)$ is:
$h(t)=h_{f}-0.5\frac{t}{T}$ (3)
  • For the ideal case: $T\to\infty$ and $h=h_{f}$
D → A: Thermalization with the hot bath.
Once the field returns to $h_{i}$ is kepet fixed, and the working substance is brought into contact with the hot bath until it returns to its initial thermal state at temperature $T_{h}$.
When analyzing the case of an infinitely slow cycle, i.e., the limit $T\to\infty$, the regime is typically referred to as adiabatic. In this regime, the unitary evolution is slow enough to satisfy the adiabatic theorem, preventing transitions between the instantaneous eigenstates of the Hamiltonian. We will calculate the following quantities based on the Appendix A [47, 48, 49]. The efficiency of the cycle can thus be expressed as:
@@
For an ideal Otto cycle, strictly speaking, perfect thermalization and perfect quantum adiabatic sweeps are ideally obtained in the limits $T\to\infty$, can be written as follows:
@@
where

### v2

@@ I Introduction
Thermodynamics is a remarkable theory in physics, primarily aimed at studying heat [1]. It has been instrumental in the design of internal combustion engines and even in explaining the behavior of black holes. Technological progress increasingly requires the miniaturization of components [2], pushing them to a regime where quantum effects are no longer negligible. The necessity of understanding the underlying mechanisms of heat conversion into useful work has led researchers to formulate the three laws of thermodynamics. More recently, with the rise of quantum technologies and the miniaturization of devices that exchange heat and work on the nanoscale, it has become increasingly relevant to investigate these mechanisms within the framework of quantum mechanics [3]. However, while classical thermodynamics is a well-established theory, its extension to the quantum domain presents conceptual challenges and remains an active area of research. Pioneering efforts to generalize the concepts of heat and work for quantum systems date back to the 1980s [4]. Only recently has a renewed interest in quantum thermal machines spurred prolific scientific advancements, starting with the groundbreaking proposal of the maser [5] as the first example of a quantum machine.
@@
A quantum thermal machine is broadly defined as a device made of quantum systems capable of performing work by undergoing a suitable thermodynamic cycle. Several studies have proposed potential implementations of few-body thermal machines [6] based on the Carnot [8, 7], Otto [9, 15], Stirling [16, 17], Brayton [18], and Diesel cycles [19]. Some of these realizations have been achieved using various quantum systems, including infinite potential wells [20], nuclear magnetic resonance (NMR) [21], ultra-cold atoms [22]. However, the role played by many-body interactions in the thermodynamic performance of quantum machines [23, 24, 25] is not yet fully understood. Determining whether quantum machines with many-body interactions can outperform classical ones, and under what conditions this occurs, is currently at the forefront of debates in quantum thermodynamics. Long-range interacting systems [26, 28] offer promising prospects for quantum technological applications due to their resilience to external disturbances [29]. This stability enables control over the impact of dynamically generated excitation, thereby mitigating their detrimental effects. Specifically, the dynamical phase transition [30, 31] is clear when $\alpha<1$. But when $\alpha>1$, the critical point grows into a chaotic crossover region where the dynamics and asymptotic state are very dependent on the system’s parameters.A promising direction for exploring quantum advantage lies in using many-body quantum systems as the working substance in thermodynamic engines [34, 35]. Although this approach is hindered by technical challenges, both analytically and numerically, some preliminary results have been obtained [36, 37, 38]. It has been shown that quantum criticality can play a crucial role in enhancing engine performance [39, 40, 41]. However, the exact role that many-body interactions play in the thermodynamic efficiency of quantum engines remains poorly understood.
The study of topological phases in systems with long-range (LR) interactions represents a dynamic research frontier, bolstered by recent experimental proposals. Notably, chains with non-local hopping and pairing are promising candidates for realization [42] in solid-state systems,an example of the rigidity of long-range interacting platforms against external drivings, is the possibility for such systems to host Floquet time crystal phases [43, 44, 45]. Theoretical insights affirm that symmetry-protected topological order can persist under LR interactions [46] in specific systems. However, the generalizability of these findings to broader symmetry-protected topological phases [45, 46, 47, 48, 49, 50] with LR interactions remains an open question, necessitating further investigation. These systems give the possibility to study a rich variety of phenomena which contrast the non-equilibrium dynamics of systems with long-range interactions to that of short-range ones, such as the violation of Lieb-Robinson bounds and anomalous propagation of information [51, 52, 53, 54, 55].
In this paper, we contribute to this burgeoning field by presenting an in-depth study of the quantum Otto engine based on the Ising model (see Fig.2), focusing particularly on the potential influence of quantum criticality on the engine’s thermodynamic performance.Our analysis also reveals that the quantum Otto engine based on the Ising model can function either as a heat engine—extracting work by absorbing heat from the cold reservoir—or as a refrigerator—using input work to transfer heat from the cold to the hot reservoir—depending on the specific parameters governing the thermodynamic cycle. While the operational regime is indeed sensitive to these parameters, one can always identify a suitable parameter range in which each mode of operation is achieved with high reliability. This study rigorously explores a quantum thermodynamic system governed by the Otto cycle, highlighting its quantum features distinct from classical counterparts [65]. We first outline the system and cycle, examining the impact of parameters such as system size $N$, initial field $h_{i}$, interaction exponent $\alpha$, and inverse temperatures $\beta_{c}$, $\beta_{h}$, on operational modes, including transitions between refrigerator and engine regimes. We then assess the quantum refrigerator and thermal engine, analyzing how $N$ and thermal correlations influence refrigeration capacity $Q_{c}/N$, work output $W/N$, and efficiencies $\eta_{R}$, $\eta$. With fixed adiabatic compression time, we investigate idealized machines with infinitely slow phases, quantifying work per spin, efficiency, and stability via the scaling factor $\Pi_{R}/N$ or $\Pi/N$ [65]. Additionally, we probe correlations for $\alpha\in[0.01,1.3]$ at $N=100$, evaluating efficiency variations with $Q_{c}/N$, $W/N$, and $h_{i}$, and their distribution across chain sizes [30, 31]. Lastly, we examine the scaling factor proportionality for $N\in[10,100]$, elucidating critical behaviors near phase transitions.
Figure 1: Dynamical phase diagram for the long-range Ising model ($J=1$), obtained via CMFT with $\ell=4$, following a quench of the transverse field from $h_{\text{in}}=0$ to $h$. The plot shows the time-averaged longitudinal magnetization (up to a time $T=1000$) as a function of $\alpha$ and $h$. Note that, as soon as $\alpha>1$ (indicated by the dashed line), the dynamical phase transition evolves into a critical region, where the asymptotic magnetization becomes highly sensitive to the system parameters. The blue square highlights.
@@ II.1 Many-body quantum Ising model
ensures that the energy density remains intensive for $\alpha\leq 1$. For $\alpha=\infty$, this model reduces to the standard transverse field Ising model (TFIM), which can be exactly solved using the Jordan-Wigner transformation [67, 66] . The long range exhibits a quantum phase transition from the ferromagnetic to the paramagnetic phase at $h_{c}=J$. This transition persists as $\alpha$ decreases [53], with the transition point shifting towards higher magnetic field values. Thus
$H=-\sum_{j=1}^{N}\sum_{r=1}^{N/2-1}\left[t_{r}c^{\dagger}_{j+r}c_{j}+\Delta_{r}c^{\dagger}_{j+r}c^{\dagger}_{j}+\text{h.c.}\right]-h\sum_{j=1}^{N}\left[1-2c^{\dagger}_{j}c_{j}\right],$ (3)
where $c^{\dagger}_{j}$ and $c_{j}$ are creation and annihilation operators for fermions at site $j$, while $t_{r}$ and $\Delta_{r}$ are the hopping and pairing amplitudes, respectively. We choose their dependence on the intersite distance $r$ according to the power laws
$t_{r}=\frac{1}{N_{\alpha_{1}}}\frac{J}{r^{\alpha_{1}}},\quad\Delta_{r}=\frac{1}{N_{\alpha_{2}}}\frac{J}{r^{\alpha_{2}}},$ (4)
with the hopping exponent $\alpha_{1}>0$, the pairing exponent $\alpha_{2}>0$, and $N_{\alpha}=\sum_{r=1}^{N/2}r^{-\alpha}$ the Kac scaling factor [31, 30], which guarantees extensivity of the energy in the case $\alpha_{i}<1$, with $i=1,2$. Hereafter, we set $J=\Delta=1$ as the energy scale and work in units of $\hbar=k_{B}=1$. This model, and with same range for pairing and hopping $\alpha_{1}=\alpha_{2}=\alpha$ [9].
Due to the translational invariant nature of the couplings, it is useful to write the Hamiltonian in terms of the momentum-space operators
$\tilde{c}_{k}=e^{-i\frac{\pi}{4}}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{ikj}c_{j},$ (5)
where $k=2\pi n/N$, with $n=-N/2+1,\dots,N/2$ (in the following we will drop the $\tilde{}$ on the $c_{k}$ unless it is ambiguous). Then we obtain
$H=\sum_{k}\left[(h-t_{k})\left(c^{\dagger}_{k}c_{k}-c_{-k}c^{\dagger}_{-k}\right)+\Delta_{k}\left(c^{\dagger}_{k}c^{\dagger}_{-k}+c_{-k}c_{k}\right)\right],$ (6)
where $t_{k}$ and $\Delta_{k}$ are the Fourier transforms of the hopping and pairing amplitudes, respectively, which in the thermodynamic limit may be written as
$t_{k}$ $=\frac{\mathrm{Re}\left[\text{Li}_{\alpha_{1}}\left(e^{ik}\right)\right]}{\zeta(\alpha_{1})},$ (7)
$\Delta_{k}$ $=\frac{\mathrm{Im}\left[\text{Li}_{\alpha_{1}}\left(e^{ik}\right)\right]}{\zeta(\alpha_{1})},$ (8)
where $\text{Li}_{\alpha}(z)$ denotes the polylogarithm and $\zeta(\alpha)$ is the Riemann zeta function. We notice that the Hamiltonian in the Fourier space can be decomposed into the sum of single mode Hamiltonians, introducing $\Psi_{k}=(c_{k},c^{\dagger}_{-k})^{T}$
$H=\sum_{k}\Psi^{\dagger}_{k}H_{k}\Psi_{k},$ (9)
$H_{k}=(h-t_{k})\sigma^{z}_{k}+\Delta_{k}\sigma^{x}_{k},$ (10)
[…]
$u_{k}=\cos\frac{\theta_{k}}{2},\quad v_{k}=\sin\frac{\theta_{k}}{2},$ (11)
where $\theta_{k}=\arctan\left[\Delta_{k}/(h-t_{k})\right]$, to obtain
$H=\sum_{k}\omega_{k}(h)\left(\gamma^{\dagger}_{k}\gamma_{k}-\frac{1}{2}\right),$ (12)
with the spectrum
$\omega_{k}(h)=2\sqrt{(h-t_{k})^{2}+\Delta_{k}^{2}}.$ (13)
@@
At the opposite extreme, $\alpha=0$ , represents a fully connected regime, which can be analytically treated for its static and dynamic properties. In this case, the model shows long-range ferromagnetic order at low finite temperatures. This configuration is particularly interesting for $\alpha<2$, revealing exotic phenomena such as prethermalization [26, 27], dynamical phase transitions [31, 30], and dynamical confinement [32, 33].
@@
This model is strongly influenced by the parameter $\alpha$. For instance, for $\alpha<2$, long-range ferromagnetic order is observed at finite temperatures [70]. The critical magnetic field is estimated using the Cluster approach and Mean-Field Theory (CMFT) [71, 72, 73] to explore the phase diagram of an Ising spin chain with long-range interactions. This method captures short-range correlations while treating interactions between clusters at a mean-field level, revealing the emergence of a chaotic region for specific values of the parameter $\alpha$. In our study, we avoid mean-field or cluster decomposition approximations and spin interaction is uniform. Nevertheless, we observe significant concordance with the results obtained via CMFT, especially concerning static phase transitions, it is possible to extract a qualitative behavior of the quantum critical point as a function of $\alpha$. Interpolating the data, we find
@@
The nonequilibrium dynamics following a quench may be probed through the time-averaged longitudinal magnetization, defined as $m_{z}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\langle\sigma_{z}(t)\rangle\,dt$, serving as the order parameter. Initiating from a fully polarized state, the system undergoes a dynamical quantum phase transition at a critical field $h_{c}=J$. In the thermodynamic limit Fig.1, for $0<\alpha<1$, the long-range Ising chain exhibits behavior analogous to the Lipkin-Meshkov-Glick model. However, as the power-law exponent surpasses the critical threshold $\alpha=1$, fluctuations become increasingly pivotal, profoundly shaping the collective dynamics of the system.
Since our method is essentially a generalized mean field theory, we will use it when the latter describes the equilibrium transition ($\alpha<5/3$). Despite this, it must be stressed that CMFT is exact for large $\ell$. We will use a numerical diagonalization of the Hamiltonian in the disordered case (Consult appendix  A and Bfor more details) .similar to [15], where we will assume $J_{ij}=\frac{1}{K(\alpha)}\frac{J}{|i-j|}$, to compute $\omega_{k}(h)$ for more details consult Otto [66]. In conclusion, the transverse field Ising model with power-law decaying interactions offers a rich theoretical framework for exploring quantum phase transitions and complex dynamics while being directly applicable to concrete experimental setups.
@@ II.2 Cycle description
We will work on an ideal otto cycle illustrated in Fig.2 composed of two infinitely slow adiabatic phases and two thermalization phases without any dissipation. After that, we will examine the influence of the thermal correlations of the $N$ particles on the efficiency of the cycle. For the dynamic and ideal case, we have:
  • (a)
    $A\to B$: Adiabatic field increase;
    The transverse field is linearly ramped up from $h_{i}$ to $h_{f}$ over time, while the working substance remains isolated from the thermal baths.
    $h(t)=h_{i}+vt,\quad t\in\left[0,\frac{h_{f}-h_{i}}{v}\right]$
  • (b)
    $B\to C$: Thermalization with the cold bath;
    The Hamiltonian $\hat{H}_{\text{sys}}(t_{f})$ is held constant, and the system is coupled to the cold bath until the working substance reaches the thermal state of $\hat{H}_{\text{sys}}(t_{f})$ at temperature $T_{c}$.
  • (c)
    $C\to D$: Adiabatic field decrease;
    The transverse field is linearly reduced from $h_{f}$ back to $h_{i}$ at the same rate as in step (a), with the working substance decoupled from the baths.
    $h(t)=h_{f}-vt,\quad t\in\left[0,\frac{h_{f}-h_{i}}{v}\right]$
  • (d)
    $D\to A$: Thermalization with the hot bath;
    The Hamiltonian $\hat{H}_{\text{sys}}(t_{i})$ is maintained, and the system is connected to the hot bath until the working substance returns to its initial thermal state of $\hat{H}_{\text{sys}}(t_{i})$ at temperature $T_{h}$.
quantities based on the Appendix A [74, 75]. The efficiency of the cycle can thus be expressed as:
@@
In what follows, we use $Q_{c}(h)$ to represent the heat exchanged during thermalization with the cold (or hot) bath, adopting the convention that $Q_{c}(h)>0$ when the system absorbs heat from the reservoir. If $\rho_{\alpha}$ is the density matrix of the system at points $\alpha=A,B,C,D$, then:
$Q_{c}$ $=\langle\hat{H}(t_{f})\rangle_{\rho_{C}}-\langle\hat{H}(t_{f})\rangle_{\rho_{B}},$ (15)
$Q_{h}$ $=\langle\hat{H}(t_{i})\rangle_{\rho_{A}}-\langle\hat{H}(t_{i})\rangle_{\rho_{D}}.$ (16)
It describes a one-dimensional long-range Ising system of $N$ quantum spins interacting with a ferromagnetic coupling strength $J>0$, in the presence of magnetic field $h$. Hereafter, we set $J=1.0$ as the energy scale and work in units where $\hbar=k_{B}=1$.The long-range interaction between spins is dictated by the strength parameter $\alpha$ (we set $h_{i}<h_{f}$ and $h_{f}-h_{i}=0.5$). When analyzing the case of an infinitely slow cycle, i.e., the limit $T\to\infty$, the regime is typically referred to as adiabatic. In this regime, the unitary evolution is slow enough to satisfy the adiabatic theorem, preventing transitions between the instantaneous eigenstates of the Hamiltonian. The nonlocal master equation chosen to model the system–environment interaction [56]. Let us consider, for a while, the more general configuration of $N_{B}$ independent thermal reservoirs at temperature $T_{n}$, with $n\in\{1,\dots,N_{B}\}$ indices labeling the bath. The Hamiltonian describing this setup reads
$\hat{H}_{\text{env}}=\sum_{n=1}^{N_{B}}\int dk\,\varepsilon_{n}(k)\hat{c}_{n}^{\dagger}(k)\hat{c}_{n}(k),$
with $\varepsilon_{n}(k)\geq 0$ and $\hat{c}_{n}(k)$, $\hat{c}_{n}^{\dagger}(k)$ being fermionic annihilation and creation operators. The $N_{B}$ baths are independent, therefore the corresponding reduced density operator of the full environment assumes the factorized form
$\rho_{\text{env}}=\bigotimes_{n=1}^{N_{B}}\rho_{\text{bath}}^{(n)},$
where $\rho_{\text{bath}}^{(n)}$ is the thermal density matrix describing the $n$th fermionic bath at temperature $T_{n}$.
[…]
Tracing out all the environmental degrees of freedom and imposing the Born–Markov approximation for the baths, it is possible to derive a microscopic Lindblad master equation , in the energy eigenbasis, for the reduced density matrix of the system described by the Hamiltonian :
$\frac{d}{dt}\rho_{\text{sys}}(t)=-i[\hat{H}_{\text{sys}},\rho_{\text{sys}}]+D[\rho_{\text{sys}}],$
$D[\rho_{\text{sys}}]=\sum_{n,k}\gamma_{nk}$ $\bigg{[}(1-f_{n}(\omega_{k}))\left(\frac{1}{2}\hat{b}_{k}\rho_{\text{sys}}\hat{b}_{k}^{\dagger}-\frac{1}{2}\{\hat{b}_{k}^{\dagger}\hat{b}_{k},\rho_{\text{sys}}\}\right)$ (17)
$+f_{n}(\omega_{k})\left(\frac{1}{2}\hat{b}_{k}^{\dagger}\rho_{\text{sys}}\hat{b}_{k}-\frac{1}{2}\{\hat{b}_{k}\hat{b}_{k}^{\dagger},\rho_{\text{sys}}\}\right)\bigg{]}.$
where the $\{\hat{b}_{k},\hat{b}_{k}^{\dagger}\}$ jump operators are the fermionic Bogoliubov quasiparticles which diagonalize the model , these operators are local in the energy eigenbasis and thus nonlocal in the sites, giving rise to a global master equation [56]. Moreover,
$f_{n}(\omega_{k})=\frac{1}{1+e^{\omega_{k}/T_{n}}},$
Under the assumption of no degeneracies in the spectrum (as turns out to be the case for the Ising chain with open boundary conditions), it can be used to obtain an analytic expression for the time evolution of the correlation functions. In particular, defining
$\tilde{f}_{k}=\frac{\sum_{n}\gamma_{nk}f_{n}(\omega_{k})}{\sum_{n}\gamma_{nk}},$
we have
$\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle_{t}=\tilde{f}_{k}\left(1-e^{-2\sum_{n}\gamma_{nk}t}\right)+\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle_{0}e^{-2\sum_{n}\gamma_{nk}t}.$
Then, all the thermodynamic quantities can be easily computed using equation (20) or, equivalently, directly computing the partition function $Z$. In particular, the internal energy reads:
$\langle E_{i}\rangle=\sum_{k}\omega_{k,i}\left(\langle b_{k}^{\dagger}b_{k}\rangle-\frac{1}{2}\right).$ (18)
For an ideal Otto cycle, in order to avoid any complications arising from the irreversibility inherent to quantum friction , we shall consider perfectly adiabatic strokes by assuming an infinitely slow transformation time $t\to\infty$ , This regime is usually referred to as adiabatic, since the unitary evolution is sufficiently slow for the adiabatic theorem to hold, preventing transitions between the instantaneous eigenstates of the Hamiltonian , which leads us back to the Fermi–Dirac distribution;
$\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle_{t\to\infty}=\tilde{f}_{k}.$
The thermodynamic quantities are calculated as follows:
@@
Figure 5: The influence of long-range interactions on the work and the quantum refrigerator’s performance per spin as a function of $h_{i}$. Here NN is fixed at 100, while $\alpha$ is varied from 0 to $1.3$.
Figure 6: In the figure, NN is fixed at 100, and $\alpha$ is varied from 0 to 1.3 to observe the role of long-range interactions on the efficiency $\eta$ of the quantum heat engine and the coefficient of performance $\eta_{R}$ of the refrigerator as a function of $h_{i}$.